compute the factor group (𝑍2 Γ— 𝑍4)/< (1,2) >

1
Expert's answer
2022-03-28T16:31:31-0400

"We\,\,have:\\\left( 0,0 \right) +\left( 1,2 \right) =\left( 1,2 \right) ,\left( 1,2 \right) +\left( 1,2 \right) =\left( 0,0 \right) \\The\,\,first\,\,class: \left( 0,0 \right) ,\left( 1,2 \right) \\Take\,\,element\,\,which\,\,was\,\,not\,\,used:\\\left( 0,1 \right) +\left( 1,2 \right) =\left( 1,3 \right) ,\left( 1,3 \right) +\left( 1,2 \right) =\left( 0,1 \right) \\The\,\,\sec ond\,\,class: \left( 0,1 \right) ,\left( 1,3 \right) \\Take\,\,element\,\,which\,\,was\,\,not\,\,used:\\\left( 1,0 \right) +\left( 1,2 \right) =\left( 0,2 \right) ,\left( 0,2 \right) +\left( 1,2 \right) =\left( 1,0 \right) \\The\,\,third\,\,class: \left( 1,0 \right) ,\left( 0,2 \right) \\Take\,\,element\,\,which\,\,was\,\,not\,\,used:\\\left( 1,1 \right) +\left( 1,2 \right) =\left( 0,3 \right) ,\left( 0,3 \right) +\left( 1,2 \right) =\left( 1,1 \right) \\The\,\,forth\,\,class: \left( 1,1 \right) ,\left( 0,3 \right) \\There\,\,are\,\,no\,\,more\,\,elements.\\The\,\,classes\,\,of\,\,elements:\\\left( 0,0 \right) ,\left( 1,2 \right) \\\left( 0,1 \right) ,\left( 1,3 \right) \\\left( 1,0 \right) ,\left( 0,2 \right) \\\left( 1,1 \right) ,\left( 0,3 \right)"